Lea now has $2,000. How much would she have after 2 years if she leaves it invested at 2% with annual compounding?


1
Expert's answer
2022-01-25T17:11:35-0500

Where 'P' refers to the principal amount, 'r' refers to the interest rate, 'n' refers to the time (in years) for which principal is borrowed and 'm' refers to the number of compounding periods.

"Compound\space interest=p-[p\times (1+ \frac{r}{m})^{n\times m}]"

"=2000\times (1+\frac{0.02}{1})^{2\times 1}\\=\$2080.80"

Thus, the amount after 2 years will be $2,080.80